Two trains travel toward each other from City A and City B, which are 400 miles apart, and they are traveling at the rates of 110 mph and 90 mph. They started at the same time and reach to City C at the same time.
a) In how many hours the trains will reach to City C between City A and City B?
b) What distance did the faster train travel to reach to City C?

Answer by lwsshak3(6465)

Two trains travel toward each other from City A and City B, which are 400 miles apart, and they are traveling at the rates of 110 mph and 90 mph. They started at the same time and reach to City C at the same time.
a) In how many hours the trains will reach to City C between City A and City B?
b) What distance did the faster train travel to reach to City C?
..
let x=distance 110 mph train traveled to city C
400-x distance 90 mph train traveled to city C
Travel time=distance/speed (equal for both trains)
x/110=(400-x)/90
90x=110*400-110x
200x=110*400
x=110*400/200=220 mi (distance traveled by first train
400-x=400-220=180 mi (distance traveled by second train
Travel time = 220/110=180/90=2 hrs
Ans:
a. The trains will reach city C in two hours
b. The faster,110 mph, train traveled a distance of 220 miles to reach city C